Given functions f and g, determine the domain of f+g, f(x)=4x/x-8, g(x)=5/x+2
Answer by nerdybill(7384)

domain of f+g, f(x)=4x/x-8, g(x)=5/x+2
.
It's all real numbers except if the denominators go to zero.
So, the exceptions are:
x-8=0
x = 8
and
x+2=0
x = -2
.
Domain (in interval notation):
(-oo, -2) U (-2, 8) U (8, oo)
where oo stands for infinity